OpenStudy (anonymous):

ln^(x^2)=((ln^x)^2) * (1/(ln^3))

OpenStudy (anonymous):

wait, let me draw it!

OpenStudy (jdoe0001):

\( \Huge ln^{x^2}=(ln^x)^2 \times \cfrac{1}{ln^3}\ \ \ ?\)

OpenStudy (anonymous):

Yes! that's the equation! sorry it was so confusing!

OpenStudy (jdoe0001):

.... aren't yo missing a base value there?

OpenStudy (anonymous):

isn't the natural log a value by its self?

OpenStudy (anonymous):

natural log or e = 2.71 i think

OpenStudy (anonymous):

the left side ln(?)^(x^2)

OpenStudy (jdoe0001):

right, I meant, the values are missing, lemme rewrite it some

OpenStudy (jdoe0001):

\(\huge ln\square^{x^2}=(ln\square^x)^2 \times \cfrac{1}{ln\square^3}\)

OpenStudy (anonymous):